y=sin-1(√(x-ax)-√(a-ax)) please find dy/dx of question

Grade:12

1 Answers

Nandana
6 years ago
hi ,
      take,
       y =sin-1(√(x-ax)-√(a-ax))
       sin y = √(x-ax)-√(a-ax)
          d/dx (sin-1 x) = 1/√(1-x2 )
         so ,
         d/dx (sin-1(√(x-ax)-√(a-ax))) = 1/√(1-(√(x-ax)-√(a-ax)2) * [-a/2 (1+lnx)*√(x-ax) +(a/2*lna* √(a-ax)]
                                                       = -½ a [(1+lnx)√(x-ax)-(lna √(a-ax )] /√(1-(√(x-ax)-√(a-ax)2)
              I think no further simplification possible !